1.     Assume a market has the following demand and supply functions:

QD = 28-3P

QS = 2P-12

Where P is the price

a)     Determine algebraically the equilibrium price and quantity.

b)     Plot the demand and supply curves and confirm your answer.

c)     Suppose supply shifts to Qs =2P – 2 with no change no change in demand. Determine the new equilibrium price and quantity.

d)    Although the supply curve shifts to the right by 10, the quantity exchanged does not rise by 10, explain why the increase in quantity is smaller.



1
Expert's answer
2021-09-06T07:07:01-0400

a)

Given

"Qd=28-3p\\Qs=2p-12"

At equilibrium Qs=Qd

"28-3p=2p-12\\28+12=2p+3p\\40=5p\\p=\frac{40}{5}\\p=8"


"Qs=2p-12\\=2(8)-12\\=16-12\\=4"


b)



c)

Given

"Qd=28-3p\\Qs=2p-2"

At equilibrium Qs=Qd

"28-3p=2p-2\\28+2=2p+3p\\30=5p\\p=\frac{30}{5}\\p=6"


"Qs=2p-2\\=2(6)-12\\=12-2\\=10"


d) Increase in supply reduce equilibrium price however it does not mean the demand will increase as there are other factors that affect demand thereby increase in quantity is smaller.